OpenStudy (republican31):

Explanation of: Find the solution set, given the replacement set. y = 1 – 3x {(1, 0), (0, 1), (1, -2), (-1, 4)} And the answer is: {(0, 1), (1, -2), (-1, 4)} I don't understand this at all. I don't remember anything about replacement sets in the lesson.

OpenStudy (anonymous):

it means a new number sets that, when plugged into the equations will still continue to be valid

OpenStudy (anonymous):

lets try all the options to show you what it means. now we try ((1,0) for\[y=1-3x\]\[0=1-3(1)\]that doesnt look valid right?

OpenStudy (anonymous):

for 2nd option (0,1)\[1=1-3(0)\] this one is valid

OpenStudy (anonymous):

for 3rd option (1,-2)\[-2=1-3(1)\] this is also valid

OpenStudy (anonymous):

now the last option (-1,4)\[4=1-3(-1)\]this one is valid too

OpenStudy (anonymous):

you can understand all that i hope

OpenStudy (republican31):

Yes thank you! So then the solutions listed are only the ones that were valid? I understand now!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yw

OpenStudy (republican31):

This was a practice problem in the lesson before the test. Now I'm ready for the test :D. Thank you so much!

OpenStudy (anonymous):

np, i didnt know about this replacement set thing until i reread the question and feel something was off lol

OpenStudy (anonymous):

good luck with your test